Consider four beakers labeled A, B, C, and D, each containing an aqueous solution and a solid piece of metal. Identity the beakers in which a chemical reaction will occur and those in which no reaction will occur.

  1. Mn(s)+CaCl2(aq)                            [ Select ]                       ["Reaction", "No reaction"]      
  2. KCl(aq)+Fe(s)                            [ Select ]                       ["Reaction", "No reaction"]      
  3. AgCl(aq)+Cu(s)                            [ Select ]                       ["Reaction", "No reaction"]      
  4. Cr(s)+HCl(aq)                            [ Select ]                       ["Reaction", "No reaction"]
  5.      

Calculate the standard free-energy change at 25 C for the following reaction:

Mg2+(aq) + Zn(s) → Mg(s) + Zn2+(aq)

Express your answer to three significant figures and in units of kJ/mol.

Which of the following redox reactions do you expect to occur spontaneously in the forward direction under standard conditions?

Check all that apply.

Group of answer choices

Ni(s)+Zn2+(aq)→Ni2+(aq)+Zn(s)

Pb(s)+Mn2+(aq)→Pb2+(aq)+Mn(s)

Ni(s)+Pb2+(aq)→Ni2+(aq)+Pb(s)

Cr(s)+3Ag+(aq)→Cr3+(aq)+3Ag(s)

None of the above
Answer #1

Solution:

1) Mn (s) + CaCl2 (aq) = No reaction

Mn can not displace Cl2 from aqueous solution of CaCl2 becuse Mn is weaker reducing agent than Mn.

2) KCl (aq) + Fe (s) = No reaction

Fe can not be able to displace Cl from KCl because Fe is weaker reducing agent than K.

3) 2AgCl (s) + Cu (s) = CuCl2 (aq) + 2Ag (s)

The reaction is single displacement reaction.It is feasible because Cu is stronger reducing agent than Ag.

4) 2Cr (s) + 6HCl (aq) = 2CrCl3 (aq) + 3H2 (g)

This is an example of single displacement reaction. Feasiblity of reaction is due to stronger reducing capacity of Cr than H.